OpenStudy (bloomlocke367):

A hyperbola has vertices of (+-5,0) and one focus (6,0). What is the standard-form of equation of the hyperbola?

OpenStudy (bloomlocke367):

is it:\[\frac{ x^2 }{ 25 }+\frac{ y^2 }{ 36 }=1\] or am I completely off?

OpenStudy (bloomlocke367):

@iGreen @SolomonZelman @satellite73

OpenStudy (bloomlocke367):

Can anyone help?..

OpenStudy (bloomlocke367):

@AlexandervonHumboldt2

OpenStudy (bloomlocke367):

@satellite73

ganeshie8 (ganeshie8):

for hyperbolas c^2 = a^2 + b^2 you're given c = 6, a = 5

ganeshie8 (ganeshie8):

find the value of \(b\) and plug in \[\dfrac{x^2}{a^2}\color{red}{-}\dfrac{y^2}{b^2}=1\]

OpenStudy (bloomlocke367):

ohhhh.. I think I so wrote that wrong.

OpenStudy (bloomlocke367):

\[b=\sqrt11\]

ganeshie8 (ganeshie8):

yep

OpenStudy (bloomlocke367):

so it's \[\frac{ x^2 }{ 25 }+\frac{ y^2 }{ 11 }=1\]

OpenStudy (bloomlocke367):

right?

ganeshie8 (ganeshie8):

careful about the sign in between

ganeshie8 (ganeshie8):

it should be \[\frac{ x^2 }{ 25 }\color{red}{-}\frac{ y^2 }{ 11 }=1\]

OpenStudy (bloomlocke367):

ohhh right. I keep messing that up.

OpenStudy (bloomlocke367):

why's it a minus sign and not a plus sign?

ganeshie8 (ganeshie8):

+ sign for ellipses - sign for hyperbolas there is a reason why this is so.. look up definition of hyperbola and ellipse

OpenStudy (bloomlocke367):

okay... thanks. :)

ganeshie8 (ganeshie8):

http://mathworld.wolfram.com/Hyperbola.html

ganeshie8 (ganeshie8):

lookup ellipse also http://mathworld.wolfram.com/Ellipse.html

OpenStudy (bloomlocke367):

thanks :)